# Elastic Softening in Synthetic Diamonds

### v2

@@ Abstract
This study reveals a novel phenomenon demonstrating the softening of synthetic diamonds when cooled to very low temperatures below 1 K. Herein, we argue that this elastic softening can be attributed to the effect of electric-quadrupole degrees of freedom of the dangling bonds in the neutral single-atom vacancies of carbon. We present the results of ultrasonic investigations of single-crystalline synthetic diamonds, namely type-IIa (colorless) and Ib (yellow) diamonds grown by high-pressure-high-temperature synthesis as well as type-IIa diamond grown by chemical vapor deposition. We observe a magnetic-field-insensitive softening of the elastic constant in all samples at low temperatures below 1 K. Our results strongly suggest a ppb level concentration of neutral single-atomic vacancies in all investigated diamonds. Our findings open new avenues for the quantitative determination of single neutral vacancies in non-irradiated diamonds, an important information needed for their potential application for quantum technology and next-generation semiconductor devices.
@@
Since ancient times, impurities and defects in diamonds have been considered as crucial criteria for appraising jewelry. Recently, the development of synthesis methods for artificial diamonds is of interest not only for semiconductor industry [1], but also for the study of the quantum properties of defects and possibilities for potential applications [2, 3, 4, 5]. Particularly, quantum states of defects arising from atomic vacancies in diamonds are considered promising for application in solid-state quantum sensors in the broad range from life to materials science [6, 7, 8, 9]. Such states also pave the way from fundamental physics to quantum technology, including quantum computation and quantum communication [10, 11, 12, 13, 14, 15]. Quantum-technology studies based on diamonds have primarily utilized commercially available synthetic diamonds prepared through high-temperature-high-pressure (HPHT) or chemical vapor deposition (CVD) methods. Although numerous researchers have investigated how defect centers originating from atomic vacancies with adjacent nitrogen atoms, called NV centers, can be grown and engineered [16, 17, 18, 19], studies comparing and evaluating the defect-center states in diamonds are scarce. In particular, whether the synthetic diamonds used as substrates for quantum devices contain atomic vacancies and to which amount has not yet been explored.
During the synthesis process, the generation of single-atom vacancies on the ppm to ppb level is thermodynamically inevitable [20] and should be well controlled in quantum technologies. Previous studies have primarily reported ppm-level concentrations of neutral single vacancies V on irradiated diamond samples [21, 22, 23]. However, the details of the existence and density of such vacancies (shown in Fig. 1) in non-irradiated synthetic diamonds have not yet been elaborated. Thus, reliable determination of V concentrations at the ppb level and investigation of vacancy properties are crucial for the rational development of novel quantum technologies. Here, we report the successful determination of the neutral single vacancy concentration in the order of ppb by quantitative analysis of the elastic softening based on the quadrupolar susceptibility [20]. Herein, the elastic “softening” refers to a decrease in the elastic stiffness constant. We further demonstrate that ultrasonic measurements at very low temperatures are a robust tool for detecting and studying neutral single vacancies in single-crystalline diamond.
Diamond has a cubic crystal structure (O, , No.227) with lattice constant Å [24], as schematically shown in Fig. 1 (top). The orbitals of carbon with the outermost electron configuration covalently bond to each other and thereby form the diamond structure, resulting in one of the hardest materials on earth. Diamonds exhibit different colors depending on the amount and type of impurities in the crystal. A yellow color is indicative for dominant nitrogen impurities (type Ib), whereas blue-colored diamond indicates the presence of boron as primary impurity (type IIb). There exist two types of colorless diamonds: type Ia, which contains nitrogen atoms in dimer form, and type IIa, which contains only a few impurities. HPHT flux and CVD methods are utilized for synthesizing type-IIa crystals. Recently, the technologies for synthesizing large high-quality single-crystalline diamonds have been significantly improved.
In an earlier study, Goto and coworkers successfully detected V vacancies [25] in single-crystalline Si through ultrasound. Indeed, Goto and coworkers observed a low-temperature softening for some acoustic modes in Si at cryogenic temperatures [20, 26, 27, 28, 29, 30, 31, 32]. Despite significant advancements of crystal-growth techniques and of the electronic properties of large-diameter Si wafers in semiconductor industry, the evaluation of quantum properties at cryogenic temperatures and the neutral single-atomic vacancy concentration leading to these properties remains to be understood properly. This is also true for diamonds. The determination of residual atomic vacancies at the few ppb level, which were previously unattainable (and overlooked) optically, is critical to improve the coherence time of quantum-information devices when downsizing quantum circuits in the future. Therefore, it is crucial to investigate the properties and establish an evaluation procedure for the low vacancy concentration in diamonds for quantum applications.
@@
We used three commercially available synthetic diamond species grown by HPHT synthesis (Sample 1 and 2) and CVD (Sample 3) for our ultrasonic measurements presented here (see SM [20] for experimental details). The temperature dependence of the elastic constant at zero magnetic field is shown in Figs. 2(a) and 2(c) for single-crystalline HPHT type-IIa diamond (Sample 1). gradually hardens down to K [as shown in the inset of Fig. 2(c)], which is a general behavior of solids owing to the anharmonicity of acoustic phonons. This change is consistent with previously reported temperature dependences of the sound velocity of transverse ultrasonic waves down to K [33, 34, 35]. Notably, to the best of our knowledge, no previous work has yet studied the elastic constants below 10 K. The low-temperature region of exhibits 125 ppm softening from 1 K down to 20 mK. In analogy to the interpretation for a similar softening observed in Si, the molecular orbitals of the neutral single vacancies remain at a triple-degenerated quantum state at least down to these temperatures. These results evidence that the local cubic symmetry around a single neutral vacancy V in diamond also remains preserved. In the Sec. 10 of the SM [20], we elaborate the reasons why the effect of NV (as shown in Fig. 1), NV (with electrons removed from NV), large multi-atom vacancy V clusters [36, 37], P centers (neutral single-substitutional nitrogen) [38], and nuclear spins have no effect on the elastic softening.
The solid red line in Fig. 2(a) is the result of calculations based on the equation (for details, see SM [20]). The fit yields mK, mK, and mK. The dashed line represents the phenomenological fit for the background contribution [39], with J m, , and = 2.941652. The negative value of indicates an antiferro-type inter-vacancy interaction for the charge state of the vacancy V. The negative value of suggests quadrupolar fluctuations associated with the triple-degenerate quantum state even at the lowest temperature of 20 mK. This result further proves that the cubic site symmetry at the vacancy site is preserved in the investigated diamond material, i.e., the local distortion owing to the static Jahn-Teller effect is irrelevant for this diamond with low vacancy concentration.
From optical measurements [25, 40, 41] on irradiated diamonds [42], an (vibronic) ground state has been proposed for V [43, 44]. This ground state cannot explain the present elastic softening in . Another experiment on non-irradiated diamond using circular-polarized light recently revealed an symmetry for the ground state of the neutral single atomic vacancy [45]. Therefore, we argue that the ground state of the neutral single vacancy (V) in the non-irradiated diamond has symmetry. A consistent picture of the quantum ground state has not been unambiguously established yet, see Sec. 9 in the SM [20].
Figure 1 illustrates schematically the crystal structure of diamond (top) and the two types of dangling bonds created by atomic vacancies (middle). There are eight atoms in the unit cell, and the atomic positions of the sites are [0,0,0], [,,0], [0,,], and [,0,]. The sites are translationally shifted by [, , ] from those. Both exhibit the same site-symmetry , where the spatial inversion symmetry is broken. Adjacent atoms are located at the vertex and center of the tetrahedron and belong to the crystallographically different and sites, respectively. As the spin-orbit interaction is small in diamond, we do not consider the effect of odd-parity multipoles but only that of even-parity ones [46]. When the molecular orbitals are placed in an electrostatic potential of site symmetry, it results in an energy splitting between a non-degenerate bonding orbital, called singlet and 3-fold degenerate antibonding orbitals, called triplet , as shown schematically at the bottom of Fig. 1 (for corresponding wave functions, see Sec. 5 in the SM [20]).
Moreover, the charge state of the atomic vacancy is determined by the site symmetry of the vacancy and number of electrons in the orbital. To describe the quantum ground state of a single neutral vacancy V, we consider a quantum state in which two electrons occupy the singlet and the other two enter the triplet state. The product space in the triplet state of the group is irreducibly decomposed as , including 9 degrees of freedom. At , the electric quadrupole degrees of freedom belonging only to the and symmetry are active. The quadrupole degrees of freedom that contribute to the softening of are , , and , which exhibit symmetry. The specific matrix elements of the electric quadrupoles can be represented using a Gell-Mann matrix and generated by applying the rotation group SU(3) to the three-dimensional representation [27].
Further, in an NV center the vacancy site exhibits lower symmetry, because of a local symmetry reduction owing to the substitution of one C site by N. Therefore, the orbital splits into two singlets and one doublet [47, 48]. When the nitrogen donor in the substituted position provides the carrier for the valence -1, the NV center is negatively charged (NV). According to Hund’s first rule, the ground state adopts the highest value of total spin , which is consistent with the Pauli exclusion principle. The six electrons occupy the singlets and the and orbits of the doublet to satisfy the condition that all wavefunctions are antisymmetric. Therefore, the ground state of NV results in an state as shown Fig. 1. However, no electric quadrupole degrees of freedom exist in the product space of with . Therefore, the orbitals of NV do not contribute to the softening of .
We estimate the vacancy concentration using the relation for the acoustic mode , which was used to analyze the vacancy concentration in Si as well [27]. Here, denotes the atomic vacancy concentration, the Jahn-Teller energy, and the deformation binding energy. The softening of Czochralski (CZ)-grown Si starts below approximately 10 K, whereas that of diamond starts below about 1 K. In line with that, obtained from the present analysis for diamond is approximately one-tenth of that of Si. Our experiments yield the parameters and , but not , which is unknown for diamond. As a first try, we assume the same value as for Si, i.e., K [29], despite the difference in the dielectric constant, melting temperature, and lattice constant [27]. The estimated neutral single-atomic vacancy concentration for the present sample of type-IIa diamond is 0.2 ppb, exhibiting the same order of magnitude as that for CZ-grown Si, namely 0.11 ppb. Notably, the large difference in the elastic constants, which is seven times larger for diamond than for Si, was also considered in these calculations. and have an error of approximately 5%. Therefore, the parameter is dominant in determining the order of .
If we use the formation energy 8.373 eV of the neutral vacancy, recently estimated by ab-initio calculation [44], we obtain a vacancy concentrations of 0.5 ppb, which is close to our above estimate of 0.02 ppb. Of course, precise values are not meaningful, but we can estimate the order of magnitude depending on how the binding constant is considered. However, there is no doubt that the vacancy concentration is definitely in the sub-ppb level. It would be highly desirable to refine the unknown parameter by theory and ultrasonic experiments using diamonds with known higher vacancy concentration, which can be controlled by irradiation.
According to the theory proposed in a previous study for Si [20, 27], the deformation binding energy can be expressed by a relationship using the quadrupole-strain coupling constant . This coupling constant is proportional to the effective charge and the square of the effective radius of the vacancy orbital (see SM) [20, 49]. In case of Si, a triple-degenerate vacancy state within the band gap can be significantly extended over a large distance of 20 Å around the vacancy site [50]. In conjunction with such theoretical predictions, ultrasonic experiments on boron-doped Si have estimated the value of the strain coupling constant , corresponding to symmetry, to be approximately K [29]. Even if the atomic vacancies are diluted, the spatially extended molecular orbitals of Si atomic vacancies contribute to the reciprocal temperature dependence of the quadrupole susceptibility, which results in the elastic softening at low temperatures. Importantly, the substantial value of allows for the detection of atomic vacancies at the sub-ppb level by use of ultrasound techniques.
@@
Further, we investigated the elastic response for other diamonds grown by different methods. Figure 3 shows the temperature dependence of for three different types of diamonds. Sample 1 is the type-IIa HPHT diamond with data already shown in Fig. 2. Sample 2 is also a type-IIa diamond grown by CVD. Sample 3 is a HPHT yellow diamond of type Ib that presumably contains numerous nitrogen impurities. Sample 1 and 2 show a nearly identical low-temperature softening within the resolution of the present measurements. The data for Sample 2 are somewhat noisy because of the small sample thickness. The softening in Sample 3 (type-Ib HPHT) is only approximately 1/3 of that of the Samples 1 and 2 (type IIa). This presumably is attributed to the neutralization of atomic vacancies by nitrogen doping. It is reasonable to assume that the concentration of neutral single vacancies in Sample 3 is less than in Samples 1 and 2 (type IIa). The inset of Fig. 3 shows data up to 100 K on a logarithmic temperature axis. Notably, the softening in Samples 2 and 3 is robust to external magnetic fields up to 14.0-16.5 T, similar as found for Sample 1 [20].

### v4

@@ Abstract
This study reveals a novel phenomenon demonstrating the elastic softening of synthetic diamonds when cooled to very low temperatures below 1 K. We present the results of ultrasonic measurements of single-crystalline synthetic diamonds without irradiation, namely type-IIa (colorless) and Ib (yellow) diamonds grown by high-pressure-high-temperature synthesis as well as type-IIa diamond grown by chemical vapor deposition. We observed a divergent decrease of the elastic stiffness constant in all samples at low temperatures down to 20 mK. We argue for the existence of electric quadrupolar degrees of freedom with irreducible representation in diamond as the origin of this elastic softening. The possible ground state preserves local symmetry presumably down to absolute zero. Of the ground states of atomic vacancies in diamond proposed thus far, the only relevant one is the positively charged single atomic vacancy V+, which has not been observed in previous optical measurements in diamonds.
I Introduction
@@
Since ancient times, impurities and defects in diamonds have been considered as crucial criteria for appraising jewelry. Diamonds exhibit different colors depending on the amount and type of impurities in the crystal. A yellow color is indicative for dominant nitrogen impurities (type Ib), whereas blue-colored diamond indicates the presence of boron as primary impurity (type IIb). There exist two types of colorless diamonds: type Ia, which contains nitrogen atoms in dimer form, and type IIa, which contains only a few impurities. Recently, the development of synthesis methods for artificial diamonds is of interest not only for semiconductor industry [1], but also for the study of the quantum properties of defects and possibilities for potential applications [2, 3, 4, 5]. Particularly, quantum states of defects arising from atomic vacancies in diamonds are considered promising for application in solid-state quantum sensors in the broad range from life to materials science [6, 7, 8, 9]. Such states also pave the way from fundamental physics to quantum technology, including quantum computation and quantum communication [10, 11, 12, 13, 14, 15] Quantum-technology studies based on diamonds have primarily utilized commercially available synthetic diamonds prepared through high-temperature-high-pressure (HPHT) or chemical vapor deposition (CVD) methods.
More recently, the technologies for synthesizing large high-quality single-crystalline diamonds have been significantly improved, and the HPHT flux and CVD methods are utilized for synthesizing type IIa crystals with extremely low vacancy concentrations. On the other hand, during the synthesis process of diamond, the generation of single-atom vacancies on the ppb level is thermodynamically inevitable [16] and should be well controlled in quantum technologies. Previous studies have primarily reported ppm-level concentrations of neutral single vacancies (V0) [shown in Fig. 1(c)] as zero phonon lines of optical centers attributed to the V0 [17, 18, 19]. The positively charged vacancy (V+) has been observed on irradiated type-IIa diamond samples in the EPR [20, 21], but it was later confirmed as negatively charged vacancy (V-) [22, 23], no further experimental verification of V+ has been made since then. Although numerous researchers have investigated how defect centers originating from atomic vacancies created by radiation damage with adjacent nitrogen atoms, called NV centers, can be grown and engineered [24, 25, 26, 27], only a few studies have evaluated the quantum states of defect centers in non-irradiated diamond because of extremely low concentration of vacancies. In particular, the quantum state of single atomic vacancies in non-irradiated synthetic diamonds used as substrates for quantum devices has not yet been explored. Thus, investigations of quantum properties of vacancies at cryogenic temperatures are crucial for the rational development of novel quantum technologies. Here, we report on the decrease in the transverse-wave elastic stiffness constants, elastic softening, revealed by ultrasonic measurements below 1 K. A possible origin of this elastic softening is the existence of electrical quadrupolar degrees of freedom, which suggests the existence of a positively charged vacancy V+ or an unresolved vacancy-derived quantum ground state with sub-ppb levels.
II Experimental Details
Diamond has a cubic crystal structure (O, , No. 227) with lattice constant Å [28], as schematically shown in Fig. 1(a). The orbitals of carbon with the outermost electron configuration are covalently bound to each other and thereby form the diamond structure, resulting in one of the hardest materials on earth. We used three commercially available synthetic diamond species grown by HPHT synthesis (Sample 1 and 2) and CVD (Sample 3) for our ultrasonic measurements presented here.
Sample 1 is a 0.16 carat mono-sectorial type-IIa (colorless) HPHT diamond (New Diamond Technology), with a length of 3.043 mm along the [110] direction and a thickness of mm along the [001] direction. According to the manufacturer, the diamond contains less than 100 ppb, or occasionally, even less than 1 ppb of nitrogen impurities in the single crystal. Sample 2 is a type-IIa CVD-grown diamond single crystal (Element Six) with a low nitrogen content of less than 5 ppb and natural abundance of 13C impurity (99.95 : 0.05 = 12C : 13C). The dimension of the sample is mm3. Sample 3 is a type-Ib (yellow) HPHT diamond (Sumitomo Electric), trade name SUMICRYSTAL (UP 100c), with a length of 2.944 mm along the [001] direction and a thickness of mm along the [100] direction.
Ultrasound was generated and detected by using a pair of LiNbO3 resonance transducers of 100 m thickness (the fundamental frequency is 18 MHz), which are bonded on the polished sample surfaces with room-temperature-vulcanizing silicone. We measured the elastic constant of sample using the transverse ultrasonic wave propagating along [110] with polarization along the [001] axis, which induced the elastic strain . Similarly, in We measured samples 2 and 3 using the transverse wave propagating along [001] with polarization along the [100] axis, which induced the elastic strain . The measured sound velocity is converted to the elastic constant using the formula [29], where g/cm3 is the calculated density of diamond. We calculated the absolute value using the measured sound velocity of 12733 m/s at 1 K.
For our low-temperature ultrasonic measurements, we used two different 3He-4He dilution refrigerators (top-loading, wet-type for Samples 1 and 3, and dry-type for Sample 2) down to 20 mK and 3He refrigerators for Samples 1 and 3 down to 450 mK. Magnetic fields up to 16.5 T were generated by superconducting magnets.
III Experimental Results
@@
The temperature dependence of the elastic constant at zero magnetic field is shown in Figs. 2(a) and 2(c) for single-crystalline HPHT type-IIa diamond (Sample 1). gradually hardens down to 1 K [as shown in Fig. 2(c)], which is a general behavior of solids owing to the anharmonicity of acoustic phonons. This change is consistent with previously reported temperature dependences of the sound velocity of transverse ultrasonic waves measures down to 10 K [32, 33, 34]. To the best of our knowledge, no previous work has yet studied the elastic constants of diamond below 1 K. The low-temperature region of exhibits 125 ppm softening from 1 K down to 20 mK. These results indicate that the sites of origin causing elastic softening locally preserve the tetrahedral symmetry even at 20 mK. Figure 2(b) shows the magnetic-field dependence of at 25 mK and 2 K with shifted offset for the 2 K data. The magnetic field is applied along the [001] direction. The 125 ppm softening is almost unchanged for magnetic fields of 16.5 T, while the data at 25 mK appear to vary slightly with respect to the magnetic field.
@@
Further, we investigated the elastic response for other diamonds grown by different methods. Figure 3 shows the temperature dependence of for three different types of diamonds. The inset of Fig. 3 shows data up to 100 K on a logarithmic temperature axis. Sample 1 is the type-IIa HPHT diamond with data already shown in Fig. 2. Sample 2 is also a type-IIa diamond grown by CVD. Sample 3 is a HPHT yellow diamond of type Ib that presumably contains numerous nitrogen impurities. Sample 1 and 2 show a nearly identical low-temperature softening within the resolution of the present measurements. The data for Sample 2 are somewhat noisy because of the small sample thickness. The softening in Sample 3 (type-Ib HPHT) is only approximately 1/3 of that of the Samples 1 and 2 (type IIa). Obviously, the vacancy or impurity concentration responsible for the softening in sample 3 is lower than in samples 1 and 2 (type IIa). Our results further imply that we can eliminate nitrogen doping as source of elastic softening. Notably, the low-temperature softening shows nearly zero magnetic-field dependence up to 14.0-16.5 T (Fig. A2 in [16]). These results suggest that the softening originates from the quantum ground state.
IV Discussion
IV.1 1: Comparison with previous studies in Si
The low-temperature softening we found in diamond (as shown in Figs. 2 and 3) is reminiscent of the elastic softening found by Goto et al. in their previous ultrasound investigations on Si, where they successfully estimated ppb levels of V0 vacancy concentrations in non-doped Si crystal and ppb levels of V+ in B-doped system grown by a floating-zone method [16, 35, 36, 37, 38, 39, 40, 41]. In analogy to the interpretation for the similar softening observed in Si, it is likely that some molecular orbitals in diamond maintain a triple degenerate electronic state down to low temperatures, i.e., the Curie-type softening of transverse mode (corresponding to the electric quadrupolar susceptibility of symmetry [16]) in diamonds could also be connected to the state in the symmetry. However, from the various experiments on irradiated diamonds [42], the ground state of the neutral vacancy V0 is believed to have vibronic many-body quantum states and with a gap of 8 meV [as shown in Fig. 1(c)]. Recent ab-initio calculations demonstrate that the dynamical Jahn-Teller-distorted structure of symmetry is lower in energy than the static Jahn-Teller distorted tetragonal structure of , in agreement with experimental observations [43]. On the other hand, the excited state, which couples to a -phonon mode, will exists above at 1.67 eV above the ground state or for V0 in irradiated diamonds [30, 44, 43, 45, 46]. Thus, it is frozen in at low temperatures and does not contribute to the low-temperature elastic softening of . In any case, these low-lying vibronic quantum ground states or do not transform as irreducible representation and are unlikely directly coupled to the -symmetric phonon mode corresponding to .
| Structure | Point | Electronic | Ground | Energy | Refs. |
| Group | Config. | State | Gap (eV) | ||
| V- | 3.150 | [47] | |||
| V0 | 1.673 | [47] | |||
| V+ | 2.8***theory [22] | [48] | |||
| NV- | 1.945 | [47] | |||
| NV0 | 2.156 | [49] | |||
| NV+ | - | [50] |
Of the ground states of atomic vacancies in diamond proposed thus far, the only relevant one is the positively charged single atomic vacancy V+. Some of the electric ground states of the various vacancy centers in diamond are listed in Table 1. Since the energetically large separation of the excited level (), the experimental information of V+ available is limited. The calculations for the V+ center suggest that the electronic state is most likely the ground state of this system [as shown Fig. 1(d)], which means that symmetry is preserved on the V+ site. The theory also demonstrates that the V+ center will not necessarily be less stable than the V0 center [48].
IV.2 2: Assessment of the concentration of positive vacancies (V+)
Here, we assume that the positive vacancy V+, which has an electric quadrupole that results in the softening of , exists in diamond. This assumption is supported by the experimental fact that the softening of type-Ib diamond with relatively high nitrogen impurity concentration is smaller than that of type-IIa, as shown in Fig. 3. Indeed, N doping changes a V+ vacancy into a NV center that no longer contributes to the lattice softening. In the following chapter, we argue that the phenomenological analysis of in Si analogously holds for diamond.
The solid red line in Fig. 2(a) is the result of calculations based on the equation (for details, see section 5 in SM [16]). The fit yields mK, mK, and the Jahn-Teller energy mK. The dashed line represents the phenomenological fit for the background contribution [29], with Jm-3, Jm-3, and = 2.941652 K. The negative value of indicates antiferro-type inter-site interactions for the electric quadrupole on the V+ states. The negative value of suggests quadrupolar fluctuations associated with the degenerate quantum state even at the lowest temperature of mK. This result further proves that the cubic site symmetry at the V+ site is preserved in the investigated diamond material, i.e., the local distortion owing to the static Jahn-Teller effect is irrelevant for diamond with low vacancy/interstitial concentration. The possible physical picture envisaged here is a situation where the electronic state of is coupled to the -phonon mode, probably forming a vibronic state due to the dynamical Jahn-Teller effect. It will allow the system to preserve tetrahedral symmetry without causing static Jahn-Teller distortion, similar to the quantum tunneling state in the clathrate crystal [51]. Another example is the non-Kramers ground state system of PrMg3, which has no long-range order down to 50 mK, where a quantum mechanical hybridization of electrons and phonons is believed to be form the vibronic state. [52].
We estimate the V+ concentration using the relation for the acoustic mode , which was used as well to analyze the vacancy concentration in Si [36]. Here, denotes the atomic vacancy concentration and the deformation binding energy. The softening of Czochralski (CZ)-grown Si starts below approximately 10 K, whereas that of diamond starts below about 1 K. In line with that, obtained from the present analysis for diamond is approximately one-tenth of that of Si. Our experiments yield the parameters and , but not , which is unknown for diamond. As a first try, we assume the same value as for Si, i.e., K (19.6 eV) [38]. The estimated V+ concentration for the present sample of type-IIa diamond is ppb, exhibiting the same order of magnitude as V0 for CZ-grown Si, namely ppb. Notably, the large difference in the elastic constants, which is seven times larger for diamond than for Si, should also be considered. In that case, the estimated value of vacancy concentration decreases to about , since is inversely proportional to square of . Since and have an error of approximately 5%. Therefore, the parameter is dominant in determining the order of . When the formation energy 8.373 eV of the neutral vacancy, recently estimated by ab-initio calculations [44] is used, we obtain a vacancy concentrations of 1.14 ppb, which is still in the ppb level. Of course, precise values are not meaningful, but we can estimate the order of magnitude depending on how the binding constant is considered. In any case, there is no doubt that the V+ concentration is definitely in the sub-ppb level in diamond without irradiation damages. It would be highly desirable to refine the unknown parameter by future surface-acoustic-wave measurements using diamonds with known higher vacancy concentration, which can be controlled by irradiation and annealing.
IV.3 3: Verification of other possible causes of softening
Next, we argue that the NV- (as shown in Fig. 1), NV0 (with electrons removed from NV-), large multi-atom vacancy Vn clusters [53], P1 centers (neutral single-substitutional nitrogen) [54], and nuclear spins do not contribute to the elastic softening found in this study. In an NV center the vacancy site exhibits the lower symmetry, because of a local symmetry reduction owing to the substitution of one C site by N. Therefore, the orbital splits into two singlets and one doublet [54, 55]. When the nitrogen donor in the substituted position provides the carrier for the valence 1, the NV center is negatively charged (NV-). According to Hund’s first rule, the ground state adopts the highest value of total spin , which is consistent with the Pauli exclusion principle. The six electrons occupy the singlets and the and orbits of the doublet to satisfy the condition that all wavefunctions are antisymmetric. Therefore, the ground state of NV- results in an state as shown Fig. 1. However, no electric quadrupole degrees of freedom exist in the product space of () with . Therefore, the orbitals of NV- do not contribute to the softening of .
This is also true for the NV0 center, large multi-atom vacancy Vn clusters, and other defects involving an asymmetric arrangement of hydrogen or boron. The P1 center, a single atomic vacancy filled by a nitrogen atom, is the most common paramagnetic impurity in natural and as-grown CVD diamonds, with , and from the 14N nucleus [54]. A previous study on the P1 center has revealed that the unpaired electron of the substitutional nitrogen occupies an antibonding orbital pointing along a C-N bond direction with most of the wave function on the nitrogen atom and a nearest-neighbor carbon atom. Notably, the impurity N relaxes off-center, and the C-N bond length increases by a few percent due to the Jahn-Teller effect, which indeed lowers the local symmetry from to . Thus, similar to NV centers, the P1 center does not cause softening that persists to 20 mK. Interactions of these defects with phonons will appear via spin-orbit coupling, which is negligible in diamond [56].
Stable isotopes of C include 12C, 13C, and 14C, with natural abundances of 98.9%, 1.1%, and 1.2 % and nuclear spin quantum numbers , and 0, respectively. None of the isotopes has a nuclear quadrupole moment; hence, the elastic constants remain unaffected at zero magnetic field. The elastic constant depends hardly on magnetic fields. The change is negligibly small compared to the low-temperature softening. Thus, we conclude that the effect of 13C is small and that nuclear spins need not to be considered.
IV.4 4: Irradiation damage
Finally, we comment on irradiation damages. There seems to be an apparent discrepancy between optical experiments on irradiated samples and ultrasonic experiments on non-irradiated samples [57]. Such discrepancy has also been pointed out in past studies of atomic vacancies in FZ-grown Si single crystals and B-doped Si single crystals [38]. In the case of Si, recent calculations indicate that it is important to treat the electron-phonon interactions (quadrupole-strain interactions) and electron correlations [58]. Similarly, in previous experiments on color centers in diamond, researchers have intentionally created a high concentration of atomic vacancies by irradiation, which leads to crystallographic damages such as atomic displacements or ionization [57]. In irradiated samples, vacancy aggregates or other thermally stable defects (interstitials) are also likely to be formed. These can induce strain fields or charge imbalance that can be felt by vacancies nearby. It might be argued that the effect of such local distortions is sufficiently small that they do not contribute to the atomic vacancy’s ground states. Indeed, it is pointed out that the acoustic modes in diamond have unusually high energies at the zone boundaries, consequently stabilizing the vacancies against a tetragonal distortion [31] and dynamic distortions result [59]. A high concentration of atomic vacancies, however, can alter the low-energy vibronic state through electron-phonon interactions. Therefore, we cannot rule out the possibility that not only the V+ but also V0, V- or other unresolved vacancies or interstitials contribute to the elastic softening found in non-irradiated diamond. These possibilities could be verified in irradiated samples by comparing the frequency dependence, ultrasound-mode dependence, and by low-temperature ultrasonic experiments utilizing surface acoustic waves.
V Conclusion